# GLDesigner: Leveraging Multi-Modal LLMs as Designer for Enhanced Aesthetic Text Glyph Layouts

### v1

@@ 4 Datasets
Given that the existing text logo datasets might be inadequate for effectively fine-tuning VLMs, we embarked on creating two extensive text logo datasets: SynTextLogo and GenTextLogo. SynthTextLogo comprises logos generated from pre-defined templates crafted by skilled designers, ensuring a diverse range of styles and typographies. Meanwhile, GenTextLogo features a curated collection of high-quality, real logos from movie and commercial posters. For each sample, we offer a comprehensive set of annotations, including bounding boxes, pixel-level masks, and character text recognitions, as well as detailed layout descriptions for every character. The subsequent subsections outline our data processing pipeline, complemented by a detailed statistical analysis of the datasets in Tab. 2. We provide examples in Fig. 1 and more can be found in supplementary.
4.1 Synthesized Text Logo Layout
@@
4.2 General Text Logo Layout
@@
To enhance data diversity, aesthetics, and align with real-world design scenarios, we have curated a comprehensive text logo dataset sourced from the internet. This collection features text logos from movie posters and commercial product advertisements, showcasing titles that are intricately stylized and expertly designed, which boasts far more complex layouts compared to traditional text fonts and predefined templates. To streamline the annotation process, we exclusively retain logos with RGBA channels, leveraging the alpha channel to effortlessly derive foreground masks. For each glyph, we meticulously annotate binary masks and text characters, supplemented by the metadata of the logo image. In total, we collect 7,068 logo samples, with 6,586 for training and 482 for testing. In comparison to the previous dataset [38], our collection offers greater diversity across logo types and styles, with longer text sequences that accommodate more complex design scenarios.
4.3 Textual Layout Description
@@
In addition to providing spatial information for each glyph, we enrich our dataset with detailed textual descriptions, which enables a deeper understanding of the overall layout and the geometric relationships between glyphs, as illustrated in Fig. 3. Leveraging off-the-shelf VLMs [1, 32, 3, 7], we automatically generate detailed layout descriptions for each glyph within a text logo, alongside a comprehensive description of the logo’s overall structure and style. Next, we employ a rigorous filtering process to ensure the accuracy of the generated descriptions by cross-verifying them with LLMs [8, 32], through prompting with generated outputs and actual position coordinates, and human annotators, thus retaining only the descriptions that accurately align with the glyphs and correcting those that do not. Most importantly, we ultimately extract and reconstruct user constraints for each logo based on the generated layout descriptions, paving the way for a user-friendly, constraint-conditioned layout generation scenario.
5 Experiments
5.1 Implement Details
@@
Unless otherwise specified, we utilized the pre-trained 7B LLaVA-1.5 model [29] with CLIP-L/336 serving as the glyph visual encoder. Each input glyph image has a size of 336x336, and for these inputs, we implemented two distinct formats: square glyphs that maintain the original aspect ratio by padding to form a square, and tight glyphs that are resized directly. During training, we randomly choose the format type, which aims to enhance the model’s ability to perceive the textural details of glyphs while preserving their original aspect ratio as closely as possible, avoiding excessive distortion. Unlike the approach in [38], which does not impose such constraints, ours seeks a balance between detail accuracy and aspect ratio fidelity. Once the layout results are generated, the entire logo is rendered using the differentiable composition techniques outlined in [38].
@@
To maintain training efficiency while preserving the model’s generalization capabilities, we keep the vision encoder frozen and only fine-tune the LLM and the projection layers. The newly added early feature projection layer is zero-initialized [15, 45] to avoid disruptive noise that could potentially compromise the stability and effectiveness of the pre-trained models.
@@
We conduct training on 8 NVIDIA V100 GPUs, with a total batch size of 128. The AdamW [30] optimizer with a learning rate of 2e-5 and a weight decay of 0.01 is employed, in conjunction with a cosine annealing scheduler that includes 3% warm-up steps. The training process integrates all three datasets for 7 epochs, which completes within 12 hours.
5.2 Evaluation Metric
For a convenient comparison with prior approaches, we utilize the same evaluation metrics as [38] including Fréchet Inception Distance (FID) [13] and Inception Score [34]. These metrics objectively assess the visual quality and diversity of the rendered logos from generated layouts.
However, the aforementioned metric falls short in fully capturing the geometric aesthetics of the generated logos, as it merely assesses the similarity between the generated and original outputs. To provide a more comprehensive evaluation of both geometric and aesthetic qualities, we additionally employ Overlap IoU, Visual Balance [41], and Glyph Ratio Consistency as quantitative metrics.
The Overlap IoU metric evaluates the ratio of pixel-level overlap rather than bounding box-level intersections in document layouts [14, 46]. Visual Balance (V.B) [41] assesses the spatial equilibrium of the generated logo, which ensures that the logo’s positioning is harmoniously centered within the canvas, contributing to an aesthetically pleasing design. Glyph Ratio Consistency (Ratio) calculates the aspect ratio difference between the generated bounding box and the original glyph image, which is an essential metric to keep the original glyph’s shape unchanged.
Most importantly, to assess how well the model adheres to the specified input constraints, we sample a subset of 50 layouts from the test set and ask human annotators to verify if the generated layouts meet the given constraints. Following [41], the average percentage of these violations is referred to the Constraint Violation Ratio (denoted as ViO).
@@ 5.2 Evaluation Metric
Table 3: Quantitative results of Unconstrained Layout Generation on the TextLogo3K [38] dataset. $\dagger$ indicates using tightly-cropped glyph images without keeping the aspect ratio, which is the default setting of [38].
@@
Table 4: Quantitative results of Unconstrained Layout Generation on our GenTextLogo dataset. $\dagger$ indicates using tightly-cropped glyph images without keeping the aspect ratio.
@@
Figure 4: Qualitative comparisons of Unconstrained Layout Generation on the TextLogo3K [38] dataset. $\dagger$ indicates using tightly-cropped glyph images without keeping the aspect ratio. The samples illustrate our model’s enhanced ability to manage long text sequences, surpassing [38].
5.3 Result Comparison
@@
Unconstrained Text Logo Layout Generation. We begin by assessing the capability to produce layouts without constraints across two distinct datasets, depicted in Tab. 3 and Tab. 4. Since our model is able to process two distinct formats of glyph images (as detailed in Sec. 5.1), for testing purposes, we utilize both square-padded and tightly-cropped glyph images, where the latter is marked with $\dagger$ for identification. When evaluating, we remove the layout requirement part from the prompt template, as outlined in Tab. 1, which allows the model full creative freedom in generating layouts.
@@
In our experiments, we also incorporated two rule-based methods, labeled as rule (a) and rule (b) in Tab. 3 and Tab. 4, where glyphs are arranged (a) in a horizontal line and (b) in a horizontal or vertical line randomly. The results illustrate that our model surpasses earlier methods, excelling in both logo quality and geometric aesthetics.
@@
Impressively, our model consistently generates aesthetically pleasing layouts while preserving the original aspect ratio of glyphs, regardless of the input format. This capability offers a significant advantage in processing stylized glyphs, as illustrated in Fig. 4, underscores our model’s perceptual ability to comprehend glyphs across various formats, thereby improving its generalization when handling unseen glyph shapes.
Figure 5: Qualitative comparisons of Unconstrained Layout Generation on our GenTextLogo dataset.
@@
User-Constrained Text Logo Layout Generation. We further conducted experiments to assess the model’s ability to generate outputs based on user-defined constraints, aligning more closely with real-world scenarios. Our findings are presented through both quantitative and qualitative analyses in Tab. 5 (utilizing the ViO metric) and Fig. 6, respectively. Due to the limitations of [38], which does not accommodate input constraints, we executed a prompt-tuning process to ensure a fair comparison. The constraints are encoded using the pre-trained CLIP Text encoder [33], and then integrated into the condition encoder described in [38], serving as an additional condition to manage the constraints effectively.
@@
However, as demonstrated by the ViO metric in Tab. 5, the fine-tuned version of [38] still struggles to comprehend and adhere to user constraints, with a violation rate exceeding 50%, which indicates that the output layout is essentially generated at random. In contrast, our model, which is enhanced with the LLM, exhibits a significantly superior ability to understand and implement user-specified constraints. Moreover, the visualization results in Fig. 6 further highlight our model’s capability to produce diverse layout arrangements while rigorously adhering to the given constraints, showcasing both flexibility and precision.
Figure 6: Qualitative results of User-constrained Layout Generation on our GenTextLogo dataset.
@@
Table 5: User studies of User-constrained Layout Generation on human annotators.
5.4 Ablation Studies
We conducted extensive ablation experiments to uncover the contributions of each component, as detailed in Tab. 6. By implementing model adjustments such as Early Feature Fusion and Adaptive Average Pooling, we observed notable improvements in mitigating glyph collision reflected in IoU scores, and reduced computational requirements (+2.4% with E.F.F and -86% with A.A.P when processing an average of 1-10 glyphs). Moreover, with the inclusion of additional datasets and textural layout description annotations, we observed significant enhancements in geometric accuracy, aesthetic quality, and layout diversity.
Additionally, augmenting the glyph inputs resulted in improved aspect ratio consistency, effectively preserving the original shapes of highly stylized glyphs, which is crucial for maintaining the integrity of complex designs.. By increasing the model’s scale from 7B to 13B, we observed further performance improvements, offering an optimal balance between computational efficiency and output quality.
@@ 5.4 Ablation Studies
Table 6: Ablation studies of   model and   training data on GenTextLogo dataset. E.F.F and A.A.P stand for Early Feature Fusion and Adaptive Average Pooling in Sec. 3.2. “detail” stands for adding textural descriptions as illustrated in Tab. 1 and Fig. 3.
Figure 7: Visualization of synthesizing text logos by combing with font generation and texture interpolation [12] in the pipeline.
5.5 User Studies
@@
In addition to evaluating generative capabilities with objective benchmarks, we conducted user studies for a subjective comparison of quality. Mirroring the ViO metric discussed in Sec. 5.2, we asked participants to choose logos they perceived to be of higher quality based on personal preference. The sampled test data included ground-truth logos sourced from the TextLogo3K and GenTextLogo datasets, alongside our generated logos using identical text content but adhering to various layout constraints.
@@
We present the preference rate, defined as the average percentage of instances where participants deemed the generated logos superior to the original ones. Additionally, we provide the average Quality ratings (from 1 to 5, worst to best), as evaluated by participants based on logo quality. The findings, detailed in Tab. 5, demonstrate that our model significantly outperforms previous methods when applied to real-world scenarios, reflecting its enhanced ability to produce visually appealing and user-preferred designs.
@@
We finally showcase a comprehensive pipeline for the automatic generation of text logos, integrating both font generation and textural interpolation [12]. This process begins with the synthesis and rendering of font glyphs based on user-specified text inputs, and prompting our model to generate an optimal layout arrangement. Finally, textural interpolation is applied to enhance the logos with rich semantic textures, elevating their visual appeal. Examples of the pipeline are showcased in Fig. 7, with slight cropping applied for space efficiency.
6 Conclusion
@@
We introduce GLDesigner, a pioneering VLM-based designing framework content-aware text logo layout generation,, capable of creating aesthetic text logos by integrating multi-modal inputs with optional user constraints natural language. We propose two innovative techniques, Early Feature Fusion and Adaptive Average Pooling, designed to significantly reduce the computational load associated with processing multiple glyph images concurrently, which results in faster training and inference without compromising performance. Additionally, we have constructed two comprehensive text logo datasets, SynTextLogo and GenTextLogo, to provide a rich source of training data essential for the instruction fine-tuning phase of our VLM. Our model excels in various benchmarks and user studies, offering a stable, controllable generation of aesthetically pleasing text logos for an enhanced user experience.

### v2

@@ 4. Datasets
Recognizing that existing text logo datasets may be insufficient for effectively fine-tuning Vision-Language Models (VLMs), we developed two comprehensive text logo datasets: SynTextLogo and GenTextLogo. SynthTextLogo comprises logos generated from pre-defined templates crafted by skilled designers, ensuring a diverse range of styles and typographies. In contrast, GenTextLogo features a curated collection of high-quality real-world logos from movie and commercial posters. For each sample, we offer a comprehensive set of annotations, including bounding boxes, pixel-level masks, and character text recognitions, as well as detailed layout descriptions for every character. The subsequent subsections outline our data processing pipeline, complemented by a detailed statistical analysis of the datasets in Table 1. Examples are shown in Figure 1, with additional examples available in the supplementary materials.
4.1. Synthesized Text Logo Layout
@@
4.2. General Text Logo Layout
@@
To enhance data diversity, aesthetics, and align with real-world design scenarios, we have curated a comprehensive text logo dataset sourced from the internet. This collection features text logos from movie posters and commercial product advertisements, showcasing titles that are intricately stylized and expertly designed, which boasts far more complex layouts compared to traditional text fonts and predefined templates. To streamline the annotation process, we exclusively retain logos with RGBA channels, leveraging the alpha channel to effortlessly derive foreground masks. For each glyph, we meticulously annotate binary masks and text characters, supplemented by the metadata of the logo image. In total, we collect 7,068 logo samples, with 6,586 for training and 482 for testing. In comparison to the previous dataset (Wang et al., 2022), our collection offers greater diversity across logo types and styles, with longer text sequences that accommodate more complex design scenarios.
4.3. Textual Layout Description
@@
In addition to providing spatial information for each glyph, we enrich our dataset with detailed textual descriptions, which enables a deeper understanding of the overall layout and the geometric relationships between glyphs, as illustrated in Figure 3. Leveraging off-the-shelf VLMs (Anthropic, 2024; OpenAI, 2024; Bai et al., 2023; Chen et al., 2023b), we automatically generate detailed layout descriptions for each glyph within a text logo, alongside a comprehensive description of the logo’s overall structure and style. We then apply a rigorous filtering process to ensure the accuracy of these generated descriptions. This involves cross-verifying them with LLMs (Dubey et al., 2024; OpenAI, 2024) by prompting with generated outputs and actual position coordinates, as well as validating through human annotators. Only descriptions that accurately align with the glyphs are retained, with corrections made where necessary. Most importantly, we ultimately extract and reconstruct user constraints for each logo based on the generated layout descriptions, paving the way for a user-friendly, constraint-conditioned layout generation scenario.
Figure 3. Textual Layout Description complemented by our data processing pipeline in Section 4.3. We provide detailed textual descriptions for the overall text logo (top) and each text glyphs (bottom), with geometric relationships between glyphs.
5. Experiments
5.1. Implement Details
@@
Unless otherwise specified, we utilized the pre-trained 7B LLaVA-1.5 model (Liu et al., 2023), with CLIP-L/336 serving as the glyph visual encoder. Each input glyph image is resized to $\times 336$ 336 pixels, and for these inputs, we implemented two distinct formats: square glyphs, which maintain the original aspect ratio by padding them to form a square, and tight glyphs, which are resized directly to fit the desired dimensions. During training, we randomly select the format type for each glyph input, which helps enhance the model’s ability to perceive the intricate textural details of glyphs while preserving their original aspect ratio as closely as possible, thereby minimizing excessive distortion. In contrast to the approach in (Wang et al., 2022), which does not impose such constraints, our method seeks a balance between detail accuracy and aspect ratio fidelity. This ensures that the generated glyphs remain visually coherent and faithful to their original design. Once the layout results are generated, the complete logo is rendered using the differentiable composition techniques outlined in (Wang et al., 2022), allowing for a seamless integration of the glyphs into the final layout.
@@
To maintain training efficiency while preserving the model’s generalization capabilities, we keep the vision encoder frozen and fine-tune only the LLM and the projection layers. The newly introduced early feature projection layer is zero-initialized (Hu et al., 2021; Zhang et al., 2023b) to prevent the introduction of disruptive noise that could potentially compromise the stability of the pre-trained models.
@@
We conduct training on 8 NVIDIA V100 GPUs, with a total batch size of 128. The AdamW (Loshchilov, 2017) optimizer with a learning rate of 2e-5 and a weight decay of 0.01 is employed, in conjunction with a cosine annealing scheduler that includes 3% warm-up steps. The training process integrates all three datasets for 7 epochs, which completes within 12 hours.
@@
Table 3. Quantitative results of Unconstrained Layout Generation on the TextLogo3K (Wang et al., 2022) dataset. $\dagger$ indicates using tightly-cropped glyph images without keeping aspect ratios.
5.2. Evaluation Metric
For a convenient comparison with prior approaches, we utilize the same evaluation metrics as (Wang et al., 2022) including Fréchet Inception Distance (FID) (Heusel et al., 2017) and Inception Score (Salimans et al., 2016). These metrics objectively assess the visual quality and diversity of the rendered logos from generated layouts.
However, the aforementioned metric falls short in fully capturing the geometric aesthetics of the generated logos, as it merely assesses the similarity between the generated and original outputs. To provide a more comprehensive evaluation of both geometric and aesthetic qualities, we additionally employ Overlap IoU, Visual Balance (Yang et al., 2024), and Glyph Ratio Consistency as quantitative metrics.
The Overlap IoU metric evaluates the ratio of pixel-level overlap rather than bounding box-level intersections in document layouts (Hsu et al., 2023; Zhou et al., 2022). Visual Balance (V.B) (Yang et al., 2024) assesses the spatial equilibrium of the generated logo, which ensures that the logo’s positioning is harmoniously centered within the canvas, contributing to an aesthetically pleasing design. Glyph Ratio Consistency (Ratio) calculates the aspect ratio difference between the generated bounding box and the original glyph image, which is an essential metric to keep the original glyph’s shape unchanged.
Most importantly, to assess how well the model adheres to the specified input constraints, we sample a subset of 50 layouts from the test set and ask human annotators to verify if the generated layouts meet the given constraints. Following (Yang et al., 2024), the average percentage of these violations is referred to the Constraint Violation Ratio (denoted as ViO).
@@ 5.2. Evaluation Metric
Table 4. Quantitative results of Unconstrained Layout Generation on our GenTextLogo dataset. $\dagger$ indicates using tightly-cropped glyph images without keeping the aspect ratio.
@@
Figure 4. Qualitative comparisons of Unconstrained Layout Generation on the TextLogo3K (Wang et al., 2022) dataset. $\dagger$ indicates using tightly-cropped glyph images without keeping aspect ratios.
5.3. Result Comparison
@@
Unconstrained Text Logo Layout Generation. We begin by assessing the capability to produce layouts without constraints across two distinct datasets, depicted in Table 3 and Table 4. Since our model is able to process two distinct formats of glyph images (as detailed in Section 5.1), for testing purposes, we utilize both square-padded and tightly-cropped glyph images, where the latter is marked with $\dagger$ for identification. When evaluating, we remove the layout requirement part from the prompt template, as outlined in Table 2, which allows the model full creative freedom in generating layouts.
@@
In our experiments, we also incorporated two rule-based methods, labeled as rule (a) and rule (b) in Table 3 and Table 4, where glyphs are arranged (a) in a horizontal line and (b) in a horizontal or vertical line randomly. The results illustrate that our model surpasses earlier methods, excelling in both logo quality and geometric aesthetics.
@@
Impressively, our model consistently generates aesthetically pleasing layouts while preserving the original aspect ratio of glyphs, regardless of the input format. This capability offers a significant advantage in processing stylized glyphs, as illustrated in Figure 4, underscores our model’s perceptual ability to comprehend glyphs across various formats, thereby improving its generalization when handling unseen glyph shapes.
@@
User-Constrained Text Logo Layout Generation. We conducted additional experiments to assess the model’s ability to generate outputs based on user-defined constraints, aiming to align the generation process more closely with real-world design scenarios. Our findings are presented through both quantitative and qualitative analyses, as shown in Table 5 (utilizing the ViO metric) and Figure 6, respectively. Due to the limitations of (Wang et al., 2022), which does not accommodate input constraints, we executed a prompt-tuning process to ensure a fair comparison. The constraints are encoded using the pre-trained CLIP Text encoder (Radford et al., 2021), and then integrated into the condition encoder described in (Wang et al., 2022), serving as an additional condition to manage the constraints effectively.
@@
However, as demonstrated by the ViO metric in Table 5, the fine-tuned version of (Wang et al., 2022) still struggles to comprehend and adhere to user constraints, with a violation rate exceeding 50%. This indicates that the generated output layouts are essentially random, failing to respect the specified requirements. In contrast, our model, which is enhanced with the LLM, exhibits a significantly superior ability to understand and implement user-specified constraints. The improvements in layout generation are particularly evident in the visualization results shown in Figure 6, where our model successfully produces a wide variety of layout arrangements while rigorously adhering to the given constraints. These results highlight not only the model’s flexibility but also its precision in meeting user specifications.
Figure 5. Qualitative comparisons of Unconstrained Layout Generation on our GenTextLogo dataset. $\dagger$ indicates using tightly-cropped glyph images without keeping aspect ratios.
@@
Table 5. User studies of User-constrained Layout Generation on human annotators. * denotes performing a prompt-tuning to accommodate input constraints.
5.4. Ablation Studies
We conducted extensive ablation experiments to investigate the contributions of each model component, as summarized in Table 6. By incorporating model adjustments such as Early Feature Fusion and Adaptive Average Pooling, we observed significant improvements, including a reduction in glyph collisions, as indicated by higher Intersection over Union (IoU) scores, and a decrease in computational requirements (+2.4% with Early Feature Fusion and -86% with Adaptive Average Pooling when processing an average of 1-10 glyphs). Moreover, the inclusion of supplementary datasets and detailed textural layout description annotations led to substantial improvements in geometric accuracy, aesthetic quality, and layout diversity.
Additionally, augmenting the glyph inputs significantly improved aspect ratio consistency, which played a crucial role in preserving the original shapes of highly stylized glyphs. This enhancement is vital for maintaining the integrity of complex designs, especially when dealing with intricate or artistic fonts. By increasing the model’s scale from 7B to 13B parameters, we observed further performance improvements, offering an optimal balance between computational efficiency and output quality. pleasing results without compromising on speed or resource usage.
Figure 6. Qualitative results of User-constrained Layout Generation on our GenTextLogo dataset.
@@ 5.4. Ablation Studies
Table 6. Ablation studies of   model and   training data. E.F.F and A.A.P stand for Early Feature Fusion and Adaptive Average Pooling in Section 3.2. ”detail” stands for adding textural descriptions as illustrated in Table 2.
5.5. User Studies
@@
In addition to evaluating the generative capabilities of our model through objective benchmarks, we conducted comprehensive user studies to provide a subjective comparison of the generated logo quality. To ensure a robust evaluation of aesthetic preferences, we adopted a methodology similar to the ViO metric, as discussed in Section 5.2. Participants were presented with pairs of logos and asked to select the one they considered to be of higher quality based on their personal preferences and visual appeal. The sampled test data used in these studies included ground-truth logos sourced from the TextLogo3K (Wang et al., 2022) and GenTextLogo datasets, alongside our generated logos using identical text content but with different layout constraints applied to reflect various design scenarios.
@@
We present the preference rate, defined as the average percentage of instances in which participants considered the generated logos superior to the original ones. Additionally, we report the average quality ratings (on a scale from 1 to 5, with 1 being the worst and 5 being the best) as evaluated by participants based on logo quality. The results, summarized in Table 5, demonstrate that our model significantly outperforms previous methods when applied to real-world scenarios, reflecting its enhanced ability to produce visually appealing and user-preferred designs.
@@
We finally present a comprehensive pipeline for the automatic generation of text logos, integrating both font generation and textural interpolation (He et al., 2023). This process begins with the synthesis and rendering of font glyphs based on user-specified text inputs, followed by prompting our model to generate an optimal layout arrangement. Subsequently, textural interpolation is applied to enrich the logos with semantic textures, elevating their visual appeal. Examples of the complete pipeline are illustrated in Figure 7, with slight cropping applied for space efficiency.
Figure 7. Pipeline of synthesizing text logos by combing with font generation and texture interpolation (He et al., 2023).
6. Conclusion
@@
We introduce GLDesigner, a pioneering VLM-based designing framework content-aware text logo layout generation, capable of creating aesthetically appealing text logos by integrating multi-modal inputs with optional user constraints in natural language. We propose two innovative techniques, Early Feature Fusion and Adaptive Average Pooling, which are designed to significantly reduce the computational load associated with processing multiple glyph images concurrently, leading to faster training and inference without compromising performance. Additionally, we have constructed two comprehensive text logo datasets, SynTextLogo and GenTextLogo, which provide a rich source of training data essential for the instruction fine-tuning phase of our VLM. Our model excels across various benchmarks and user studies, offering a stable and controllable generation of visually appealing text logos, thereby enhancing the overall user experience.